# How Panel Layouts Define Manga: Insights from Visual Ablation Experiments

### v1

@@ Abstract
Today, manga has gained worldwide popularity. However, the question of how various elements of manga, such as characters, text, and panel layouts, reflect the uniqueness of a particular work, or even define it, remains an unexplored area. In this paper, we aim to quantitatively and qualitatively analyze the visual characteristics of manga works, with a particular focus on panel layout features. As a research method, we used facing page images of manga as input to train a deep learning model for predicting manga titles, examining classification accuracy to quantitatively analyze these features. Specifically, we conducted ablation studies by limiting page image information to panel frames to analyze the characteristics of panel layouts. Through a series of quantitative experiments using all 104 works, 12 genres, and 10,122 facing page images from the Manga109 dataset, as well as qualitative analysis using Grad-CAM, our study demonstrates that the uniqueness of manga works is strongly reflected in their panel layouts.
@@ Introduction
Manga, a distinctive form of Japanese comics, is characterized by its unique panel arrangements, which, along with creative character designs, guide the reader’s eye through the story in a carefully orchestrated manner. Panel layouts play a crucial role in conveying the narrative, as emphasized by Shogakukan’s “Newcomer Manga Award Manga Creator Training Course” [^1] https://shincomi.shogakukan.co.jp/training/004.html, which outlines five fundamental rules for panel layout design. Manga artists can further enhance originality by creatively adhering to or intentionally breaking these rules to suit their storytelling needs. However, while readers may feel the subtle differences of panel layouts across manga, they often find it challenging to clearly distinguish the specific layout differences between works or identify whether a work has a unique panel layout of its own. This raises an important question: To what extent do panel layouts reflect the uniqueness of a manga work, and how do they contribute to its overall identity?
@@
The results show that panel layouts possess inherent uniqueness, even in the absence of content such as characters or text. This underscores the critical role of panel layouts in distinguishing manga works and highlights their potential as a unique representation of an author’s creative intent. These findings contribute to the understanding of panel layouts as an essential and unique element in manga creation, offering new insights into the creative processes of manga authors.
@@
There have been numerous studies that use images of creative works as input to predict attributes such as the title or genre associated with those images. \@BBOPcitet\@BAP\@BBNChu & Guo (2017)\@BBCP, for instance, used a movie poster dataset and fine-tuned a pre-trained CNN to perform classification of the movie genres based on the poster images. Similarly, \@BBOPcitet\@BAP\@BBNWi et al. (2020)\@BBCP built a CNN model to predict the genre of movies from their poster images, and later visualized which parts of the image the CNN focused on for classification using Grad-CAM, verifying whether the CNN was extracting the expected features from the poster images. There are also studies on the role of panel layouts. \@BBOPcitet\@BAP\@BBNGodwin et al. (2018); Takahashi & Kondo (2023)\@BBCP demonstrated that the layout of books affects reading comprehension in children and university students, respectively. \@BBOPcitet\@BAP\@BBNTüchler et al. (2021)\@BBCP highlighted the impact of webpage layouts on user experience.
Recently, several experiments have emerged that utilize deep learning methods to study comics and manga. \@BBOPcitet\@BAP\@BBNLaubrock & Dunst (2020); Augereau et al. (2018); Sharma & Kukreja (2024)\@BBCP had introduced computer techniques used in comics research in recent years. \@BBOPcitet\@BAP\@BBNLaubrock & Dubray (2019)\@BBCP conducted classification experiments on comic titles and authors using CNNs, leveraging large-scale English comic datasets such as GNC \@BBOPcitep\@BAP\@BBN(Dunst et al., 2017)\@BBCP and Manga datasets such as Manga109 \@BBOPcitep\@BAP\@BBN(Fujimoto et al., 2016; Matsui et al., 2017)\@BBCP. Additionally, \@BBOPcitet\@BAP\@BBNDubray & Laubrock (2019)\@BBCP used the GNC English comic dataset to perform object detection experiments on speech balloons within comic pages.  \@BBOPcitet\@BAP\@BBNRigaud et al. (2021)\@BBCP also conducted experiments on detecting text within speech balloons in comics. \@BBOPcitet\@BAP\@BBNYoung-Min (2019)\@BBCP used CNNs to classify manga pages and employed feature visualization techniques to analyze the internal visual features of the model. \@BBOPcitet\@BAP\@BBNKukreja et al. (2023)\@BBCP conducted classification experiments on individual manga pages using CNNs and Random Forest.
@@
As described in the above subsection, various studies have been conducted using manga page image data. However, most of these studies aim to apply manga image data for engineering purposes, and there are few studies that attempt to analyze the features of manga images from a scientific perspective. While the classification experiments on manga page images conducted by  \@BBOPcitet\@BAP\@BBNLaubrock & Dubray (2019)\@BBCP can be considered an attempt to analyze manga features from a scientific viewpoint, their experiment did not analyze individual components, such as panel layouts, but rather limited their investigation to quantitatively examining the characteristics of works in terms of classification accuracy.
Furthermore, no one has attempted to classify manga works using deep learning models by inputting processed manga images, such as masked images or panel frames, and then qualitatively analyzing the distinct panel layout features of each work through visualizing key regions in the classification model. In \@BBOPcitet\@BAP\@BBNKukreja et al. (2023)\@BBCP, while they trained a CNN model to classify manga genres, they did not consider whether each manga possesses unique stylistic characteristics. During training, the model might have learned features unique to individual works rather than focusing solely on the genre. In  \@BBOPcitet\@BAP\@BBNKukreja et al. (2023)\@BBCP, 8 manga books were selected for classification to avoid this issue. However, the dataset was too small, and the experiments were limited to panels and pages, without further investigating features such as characters or panel layouts independently.
@@
For rectangular noisy panels, the model achieved an average classification accuracy of 67.2% with noise of (-10, 10) and 52.9% with noise of (-20, 20). In contrast, quadrilateral noisy panels resulted in slightly higher accuracies of 70.7% and 59.3%, respectively. Although classification accuracy for noisy panel frame images was lower than that for original panel frame images, the model still demonstrated a certain level of accuracy. These results indicate that while noise reduces the model’s ability to utilize panel layout features, the distinctive characteristics of panel designs across works remain sufficient for classification, even under noisy conditions.
@@ Conclusion
This study used deep learning to explore whether panel page designs in manga vary by work. Our experiments showed that even without characters and text, panel layouts exhibit inherent uniqueness, serving as a key distinguishing feature for manga. This was validated through classification tasks and supported by Grad-CAM visualizations.
Additional experiments confirmed that publishers and genres have minimal influence on panel layouts, highlighting their individuality at the work level. Even with added noise, panel layouts retained sufficient distinctiveness for classification, emphasizing their intricate role in manga creation.
These findings highlight the significance of panel layouts as a unique visual feature in manga, reflecting not only the creative diversity of authors but also their storytelling intent. Future research directions include conducting experiments with a larger dataset to verify whether classification accuracy can remain high as the number of categories increases. Additionally, due to current data limitations, the differences in page design between works by the same author have not been thoroughly explored. Further investigation into this question is planned for future studies.

### v2

@@ Abstract
Manga has gained global popularity, yet how its visual elements, such as characters, text, and panel layouts, reflect the uniqueness of individual works remains underexplored. This study investigates the contribution of panel layouts to manga identity through both quantitative and qualitative analysis. We trained a deep learning model to classify manga titles based solely on facing page images, and performed ablation experiments by removing characters and text, retaining only panel frame structures. Using 10,122 images from 104 works in 12 genres in the Manga109 dataset, we demonstrate that panel layouts alone enable high-accuracy classification. Grad-CAM visualizations further reveal that the models focus on layout features such as size, spacing, and alignment. These findings suggest that panel layouts encode work-specific stylistic patterns and support visual narrative comprehension, highlighting their role as a key component of manga’s visual identity.
@@ Introduction
Manga, a distinctive form of Japanese comics, is characterized by its unique panel arrangements, which, along with creative character designs, guide the reader’s eye through the story in a carefully orchestrated manner. Panel layouts play a crucial role in conveying the narrative, as emphasized by Shogakukan’s “Newcomer Manga Award Manga Creator Training Course” [^1] https://shincomi.shogakukan.co.jp/training/004.html, which outlines five fundamental rules for panel layout design. Manga artists can further enhance originality by creatively adhering to or intentionally breaking these rules to suit their storytelling needs. While readers may perceive subtle layout differences, it is often difficult to determine how these reflect the style of a work. This raises an important question: To what extent do panel layouts reflect the uniqueness of a manga work, and how do they contribute to its overall identity?
@@
These results indicate that panel layouts possess inherent uniqueness, even in the absence of content such as characters or text. This suggests that layout structures carry stylistic information that may reflect the unique visual identity of each work. Although our approach is based on a classification task, the broader aim is to understand how layout alone encodes work-specific style and supports visual narrative interpretation. This line of inquiry is closely related to cognitive theories of attention and sequential image comprehension.
@@
Numerous studies have used images of creative works as input for predicting associated attributes such as titles or genres. \@BBOPcitet\@BAP\@BBNChu & Guo (2017)\@BBCP fine-tuned a pretrained CNN to classify movie genres based on poster images. \@BBOPcitet\@BAP\@BBNWi et al. (2020)\@BBCP applied a similar approach and visualized the model’s focus using Grad-CAM to verify whether it extracted expected features. \@BBOPcitet\@BAP\@BBNGodwin et al. (2018); Takahashi & Kondo (2023); Tüchler et al. (2021)\@BBCP examined how layout affects comprehension or user experience, including children’s book design and web interfaces.
Several works also explored deep learning approaches in comic research. \@BBOPcitet\@BAP\@BBNLaubrock & Dunst (2020); Augereau et al. (2018); Sharma & Kukreja (2024)\@BBCP surveyed computer vision applications in comics. \@BBOPcitet\@BAP\@BBNLaubrock & Dubray (2019)\@BBCP conducted classification of comic authors and titles using CNNs on datasets like GNC \@BBOPcitep\@BAP\@BBN(Dunst et al., 2017)\@BBCP and Manga109 \@BBOPcitep\@BAP\@BBN(Fujimoto et al., 2016; Matsui et al., 2017)\@BBCP. \@BBOPcitet\@BAP\@BBNDubray & Laubrock (2019)\@BBCP performed speech balloon detection, while  \@BBOPcitet\@BAP\@BBNRigaud et al. (2021)\@BBCP worked on text segmentation inside speech balloons. \@BBOPcitet\@BAP\@BBNYoung-Min (2019)\@BBCP used CNN-based classification and feature visualization. \@BBOPcitet\@BAP\@BBNKukreja et al. (2023)\@BBCP applied CNN and Random Forest models to classify manga pages.
@@
While a growing number of studies have used manga page images, most focus on engineering applications, with limited attention to the underlying visual structure of manga. \@BBOPcitet\@BAP\@BBNLaubrock & Dubray (2019)\@BBCP attempted to analyze manga features through classification, but their method treated pages as undivided wholes, without isolating individual components like panel layout. \@BBOPcitet\@BAP\@BBNYoung-Min (2019)\@BBCP proposed a CNN-based model for author classification and visualized learned features, but their analysis focused on stylistic textures and drawing patterns, rather than structural layout.
Although previous studies have explored page-level classification, none have specifically isolated panel layouts through ablation and visualization to analyze their contribution to manga identity. Although \@BBOPcitet\@BAP\@BBNKukreja et al. (2023)\@BBCP trained a CNN to classify genres, they did not consider whether each manga has unique stylistic traits. Their dataset was small and did not evaluate features like panel layout independently. In contrast, our work focuses on isolating panel layout features through ablation and visual interpretation, to investigate their role in distinguishing manga works.
@@
Table 4: Classification accuracy on noisy panel frame-only images.
Noise Type Noise Range Accuracy
Rectangular (-10, 10) 0.67
Rectangular (-20, 20) 0.53
Quadrilateral (-10, 10) 0.71
Quadrilateral (-20, 20) 0.59
We further added noise to panel frame positions to test the model’s sensitivity to layout distortion. As shown in Table 4, the classification accuracy decreased as the noise level increased, indicating that precise panel positioning plays a meaningful role in layout-based classification. Although classification accuracy for noisy panel frame images was lower than that for original panel frame images, the model still demonstrated a certain level of accuracy. These results indicate that while noise reduces the model’s ability to utilize panel layout features, the distinctive characteristics of panel designs across works remain sufficient for classification, even under noisy conditions.
@@ Conclusion
This study used deep learning to explore whether panel page designs in manga vary by work. Our experiments showed that even without characters and text, panel layouts exhibit work-specific uniqueness, enabling reliable classification and highlighting their role as a distinguishing visual structure in manga. This was validated through classification tasks and supported by Grad-CAM visualizations.
Additional experiments confirmed that panel layouts are minimally influenced by publishers or genres, underscoring their individuality at the work level. Even with added noise, panel layouts retained sufficient distinctiveness for classification, emphasizing their intricate role in manga creation.
While our study isolates panel layouts, we acknowledge that other visual elements—such as character design, gaze direction, and speech balloon positioning—also influence manga style and storytelling. Future research could investigate the interplay between these features and layout.
These findings highlight the significance of panel layouts as a unique visual feature in manga, reflecting the stylistic diversity of individual works and their contribution to storytelling structure. Future research directions include conducting experiments with a larger dataset to verify whether classification accuracy can remain high as the number of categories increases. Additionally, due to current data limitations, the differences in page design between works by the same author have not been thoroughly explored. Further investigation into this question is planned for future studies.
Acknowledgments
This work was supported by JSPS KAKENHI Grant Number JP24K02993.